OpenStudy (anonymous):

I need help with The half-life of a certain radioactive material is 32 days. An initial amount of the material has a mass of 361 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth.

OpenStudy (anonymous):

The general form is N(t)/N(0) = exp(-kt) N(t) is the amount (mass, number) of radioactive atoms at time t. Get k by knowing one such ratio.Here, the half-life is 32 days, N(32)/N(0) = 0.5 = exp(-32 k) ln (0.5) = - 0.693 = - 32 k k = 0.0217 / day. Put this k into exp(-kt) with t = 5 days and now have fraction left. (nearly 90%) Multiply fraction left by original mass and get final mass.

OpenStudy (anonymous):

this is the one I don't understand I picked A which was wrong

OpenStudy (anonymous):

so what is the final mass

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

The equation I use for this problem is m(t)= Ce^kt \[180.5=361e ^{32k}\] \[\frac{ 180.5}{ 361 } = e ^{32k}\] \[\ln \frac{ 1 }{ 2}= \ln e ^{32k}\] \[\ln \frac{ 1 }{ 2 }= 32k\] \[\frac{ 1 }{ 32}\ln \frac{ 1 }{ 2 }= k\] m(t)= \[361e ^{1/32\ln(1/2) t}\] That's your equation. then you plug in your time t=5 to find how much material remains.

OpenStudy (anonymous):

I don't have a calculator on hand right now

OpenStudy (anonymous):

323.945